OpenStudy (anonymous):

The difference between the length of the hypotenuse and the length of the next longest side of a right triangle is 3 cm. The difference between the lengths of the two perpendicular sides is 3 cm. Find the three side lengths.

OpenStudy (anonymous):

I don't get it..

OpenStudy (turingtest):

call the hypotenuse length x:use the Pythagorean theorem:\[x^2=(x-3)^2+(x-6)^2\]and solve for x.

OpenStudy (anonymous):

why is it x-3, and x-6?

OpenStudy (anonymous):

as The difference between the lengths of the two perpendicular sides is 3 cm

OpenStudy (anonymous):

So once I get: \[(x ^{2}-6x+9)(x ^{2}-12x+36)\]

OpenStudy (anonymous):

what do I do after?

OpenStudy (anonymous):

utzz...rong u r missin +sign in between

OpenStudy (anonymous):

\[(x ^{2}-6x+9)+(x ^{2}-12x+36)=x ^{3}\]

OpenStudy (anonymous):

ohh, yup I put that now. And so I add to: 2x^2 - 18x + 45?

OpenStudy (anonymous):

imean....x^2

OpenStudy (anonymous):

equate to x^2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

so the equation becomes....... 2x^2 - 18x + 45=x^2

OpenStudy (anonymous):

hence 2x^2 - 18x + 45-x^2=0

OpenStudy (anonymous):

x^2 - 18x + 45?

OpenStudy (anonymous):

x^2 - 18x + 45=0

OpenStudy (anonymous):

=> (x-3)(x-15)=0

OpenStudy (anonymous):

hence x=3 or x=15

OpenStudy (anonymous):

got it?/

OpenStudy (turingtest):

...and since x=3 will not produce a triangle we have x=15

OpenStudy (anonymous):

yeah right...

OpenStudy (anonymous):

ohh..